# NarraBench: A Comprehensive Framework for Narrative Benchmarking

### v1

@@ 6 Conclusion
To help guide and centralize growing efforts in narrative benchmarking, we present NarraBench, the first comprehensive narratological taxonomy for benchmark developers to integrate efforts around a shared, holistic understanding of narrative communication. NarraBench distills decades of literary theory down to four fundamental narrative dimensions any NLP researcher should be aware of when engaging with narratives in any medium: story, narration, discourse (or the structure of the narrative), and situatedness (the context in which the narrative exists). We further sub-divide these dimensions into twelve primary features and fifty overall aspects. We present the full taxonomy together with example questions in Table 2.
As a first step towards the construction of the full suite of benchmarks indicated by our framework, we conducted a first-ever survey of existing narrative benchmarks to identify how contemporary benchmarking efforts align with the NarraBench taxonomy (Table 3). We identify 39 open-data benchmarks whose qualitative targets roughly match one of the fifty permutations defined in our taxonomy, assessing each benchmark’s fit according to our specified evaluation criteria. This process reveals present benchmarks satisfy approximately 27% of the full NarraBench taxonomy as indicated in Figure 1. This gap suggests there are significant opportunities available for benchmark developers to produce new tasks for testing model narrative understanding.
@@
We foresee two ways NLP practitioners can benefit from, and contribute to the NarraBench framework:
  1. 1.
    Community involvement. We encourage all interested to invest in expanding narrative benchmarking in areas not yet covered as per our initial survey (cf. Table 3). To that end we are maintaining a live spreadsheet here. The spreadsheet will allow community members to consult the current state of the art, add on additional taxonomic categories when appropriate, and to submit their own benchmarks as they are being developed. Distributing a stable web resource is one important way NarraBench will remain an evolving and relevant resource moving into the future.
  2. 2.
    Reference benchmark implementation. NarraBench is a novel attempt at reconfiguring existing benchmarks to fulfill theoretical goals, and as such can be implemented as an expandable collection of relevant benchmarks. To guide future implementations we will collect and script the most suitable benchmarks identified in our study to produce a reference implementation. This testing harness will evolve and grow as benchmarks for the missing rows in Table 3 are developed and/or submitted to the NarraBench organizers. Community members are likewise welcome to adopt our taxonomy for (re)implementation with proprietary methods and datasets according to their needs.
A growing body of research is increasingly focused on studying the value of narrative understanding and narrative generation. Now is an optimal time to begin the work of consolidation to provide LLM developers and researchers with clear assessment criteria around one of the most fundamental ways that humans relate to each other: storytelling.
@@
While NarraBench provides the first comprehensive taxonomy for narrative benchmarking, several limitations should be acknowledged. First, the framework is grounded in a particular lineage of narrative theory—what is known as the classical model—which privileges compositional and communicative dimensions of narrative (story, narration, discourse, situatedness). This excludes alternative conceptions of narrativity, such as those derived from cognitive narratology, rhetorical theory, or postclassical approaches emphasizing affect, ideology, or embodiment. As a result, NarraBench captures a specific, though widely applicable, view of what constitutes narrative understanding in computational terms.
@@
Finally, despite its explicit focus on perspectival and consensus-based evaluation, NarraBench inherits many of the broader evaluation challenges facing LLM benchmarking: the instability of LLM-as-a-judge methods, limited reproducibility of generative scoring, and uncertainty about how human interpretive diversity should be represented in gold standards. Expanding community participation, diversifying annotation pipelines, and incorporating non-deterministic scoring schemes will be essential to realizing the full potential of narrative benchmarking.
In sum, NarraBench represents a first synthesis—a structured, extensible foundation for assessing narrative understanding—while acknowledging that both its theoretical premises and empirical coverage are partial. Its value lies in providing a coherent starting point for an evolving research community to debate, test, and revise what narrative understanding should mean for artificial intelligence.
@@ Appendix B Survey Results
role Crab global holistic consensus (P) He et al. (2025)
Ditto global holistic deterministic (P) Lu et al. (2024)
attributes AustenAlike global holistic consensus Yang and Anderson (2024)
NoLiMa global holistic deterministic (C) Modarressi et al. (2025)
emotions CULEMO global holistic perspectival Belay et al. (2025)
EQ-Bench global holistic deterministic (P) Paech (2024)
ToMBench global holistic deterministic (P) Chen et al. (2024b)
EmotionBench global holistic deterministic (P) Chen et al. (2024a)
motivation OpenToM local progressive (D) deterministic (P) Xu et al. (2024)
FANTOM global holistic (P) deterministic (P) Kim et al. (2023)
MOTIVEBENCH global holistic (P) deterministic (P) Yong et al. (2025)
CHARACTERBENCH global holistic (P) deterministic (P) Zhou et al. (2024)
@@
connections PhantomWiki global holistic deterministic Gong et al. (2025)
relationship DialogBench global holistic deterministic (P) Ou et al. (2024)
@@
Plot topic RepLiQA global holistic deterministic (C) Monteiro et al. (2024)
plot FlawedFictions global progressive (H) deterministic (P) Ahuja et al. (2025)
plotline StorySumm global holistic consensus Subbiah et al. (2025)
moral MoralBench global holistic deterministic (P) Ji et al. (2024)
MORABLES global holistic deterministic (P) Marcuzzo et al. (2025)
@@
Setting setting Dataset-GSS global holistic deterministic (C) Li et al. (2025)
@@
Structure plot arc EventStoryLine global holistic (P) deterministic (C) Caselli and Vossen (2017)
Narrative-Discourse global holistic (P) deterministic (C) Tian et al. (2024)
Perspective dialogue SODA-EVAL global (L) holistic (H) deterministic Mendonça et al. (2024)
CoSER global (L) holistic (H) deterministic Wang et al. (2025d)
@@
order TRaVelER global holistic deterministic Kenneweg et al. (2024)
ToT global holistic deterministic Fatemi et al. (2024)
MAVEN-ERE global holistic deterministic Wang et al. (2022a)
TRAM global holistic deterministic Wang and Zhao (2024)
Revelation suspense ConflictBank global holistic (P) deterministic (P) Su et al. (2024)
@@
Paratext genre BookCover30 global discrete (H) deterministic (C) Iwana et al. (2017)
author VeriDark global holistic (D) deterministic Manolache et al. (2022)
REASONS global discrete deterministic Saxena et al. (2025)
date TimeAware global discrete deterministic Herel et al. (2025)
@@
Motivation intent Persuasive Pairs global holistic deterministic (P) Pauli et al. (2025)

### v3

@@ 7 Conclusion
To help guide and centralize growing efforts in narrative benchmarking, we present NarraBench, the first comprehensive narratological taxonomy for benchmark developers to integrate efforts around a shared, holistic understanding of narrative communication. NarraBench distills decades of literary theory down to four fundamental narrative dimensions: story, narration, discourse (or the structure of the narrative), and situatedness (the context in which the narrative exists). We further divide these dimensions into twelve primary features and fifty overall aspects, described in Table 4.
As a first step towards constructing our envisioned framework, we conducted a first-ever survey of existing narrative benchmarks to identify how contemporary benchmarking efforts align with the NarraBench taxonomy (Table 5). We identify 39 open-data benchmarks whose qualitative targets roughly match one of the fifty permutations defined in our taxonomy, suggesting existing benchmarks satisfy $\approx$27% of the NarraBench taxonomy as indicated in Figure 1. This gap suggests opportunities for researchers to produce new tasks.
@@
Researchers are increasingly focusing on the value of narrative understanding and narrative generation. Our testing harness will allow community members to consult the current state of the art, and contribute where possible. Now is an optimal time to begin the work of consolidation to provide LLM researchers with clear assessment criteria around one of the most fundamental ways humans relate to each other: storytelling.
@@
While NarraBench provides the first comprehensive taxonomy for narrative benchmarking, several limitations should be acknowledged. First, the framework is grounded in a particular lineage of narrative theory—what is known as the classical model—which privileges compositional and communicative dimensions of narrative (story, narration, discourse, situatedness). This excludes alternative conceptions of narrativity, such as those derived from cognitive narratology, rhetorical theory, or postclassical approaches emphasizing affect, ideology, or embodiment. As a result, NarraBench captures a specific, though widely applicable, sense of computational narrative understanding.
@@
Finally, despite its explicit focus on perspectival and consensus-based evaluation, NarraBench inherits many of the broader evaluation challenges facing LLM benchmarking: the instability of LLM-as-a-judge methods, limited reproducibility of generative scoring, and uncertainty about how human interpretive diversity should be represented in gold standards. Expanding community participation, diversifying annotation pipelines, and incorporating non-deterministic scoring schemes will be essential to properly realizing narrative benchmarking.
In sum, NarraBench represents a first synthesis—a structured, extensible foundation for assessing narrative understanding—while acknowledging that both its theoretical premises and empirical coverage are partial. Its value lies in providing a coherent starting point for an evolving research community to debate, test, and revise what narrative understanding should mean for NLP and ML.
@@ Appendix B Survey Results
role Crab global holistic consensus (P) heCrabNovelConfigurable
Ditto global holistic deterministic (P) luLargeLanguageModels2024
attributes AustenAlike global holistic consensus yangEvaluatingComputationalRepresentations2024
NoLiMa global holistic deterministic (C) modarressiNoLiMaLongContextEvaluation2025
emotions CULEMO global holistic perspectival belayCULEMOCulturalLenses2025
EQ-Bench global holistic deterministic (P) paechEQBenchEmotionalIntelligence2024
ToMBench global holistic deterministic (P) chenMBENCHBenchmarkingTheory
EmotionBench global holistic deterministic (P) chenEmotionQueenBenchmarkEvaluating2024
motivation OpenToM local progressive (D) deterministic (P) xuOpenToMComprehensiveBenchmark2024
FANTOM global holistic (P) deterministic (P) kimFANToMBenchmarkStresstesting2023
MOTIVEBENCH global holistic (P) deterministic (P) yongMotiveBenchHowFar2025
CHARACTERBENCH global holistic (P) deterministic (P) zhouCharacterBenchBenchmarkingCharacter
@@
connections PhantomWiki global holistic deterministic gongPhantomWikiOnDemandDatasets2025
relationship DialogBench global holistic deterministic (P) ouDialogBenchEvaluatingLLMs2024
@@
Plot topic RepLiQA global holistic deterministic (C) monteiroREPLIQAQuestionAnsweringDataset
plot FlawedFictions global progressive (H) deterministic (P) ahujaFindingFlawedFictions2025a
plotline StorySumm global holistic consensus subbiahSTORYSUMMEvaluatingFaithfulness2025
moral MoralBench global holistic deterministic (P) jiMoralBenchMoralEvaluation
MORABLES global holistic deterministic (P) marcuzzoMORABLESBenchmarkAssessing2025
@@
Setting setting Dataset-GSS global holistic deterministic (C) liPixelsPlacesSystematic2025
@@
Structure plot arc EventStoryLine global holistic (P) deterministic (C) caselliEventStoryLineCorpus2017
Narrative-Discourse global holistic (P) deterministic (C) tianAreLargeLanguage2024
Perspective dialogue SODA-EVAL global (L) holistic (H) deterministic mendoncaSodaEvalOpenDomainDialogue2024
CoSER global (L) holistic (H) deterministic wangCoSERCoordinatingLLMBased2025
@@
order TRaVelER global holistic deterministic kennewegBenchmarkingAbilityLarge2024
ToT global holistic deterministic fatemiTestTimeBenchmark2024
MAVEN-ERE global holistic deterministic wangMAVENEREUnifiedLargescale2022
TRAM global holistic deterministic wangTRAMBenchmarkingTemporal2024
Revelation suspense ConflictBank global holistic (P) deterministic (P) suConflictBankBenchmarkEvaluating2024
@@
Paratext genre BookCover30 global discrete (H) deterministic (C) iwanaJudgingBookIts2017
author VeriDark global holistic (D) deterministic manolacheVeriDarkLargeScaleBenchmark2022
REASONS global discrete deterministic saxenaAttributionScientificLiterature2025
date TimeAware global discrete deterministic herelTimeAwarenessLarge2025
@@
Motivation intent Persuasive Pairs global holistic deterministic (P) pauliMeasuringBenchmarkingLarge2025